You are at a bank to setup a bank account with an ATM card. The bank requires you to enter a 4-digit PIN, and each digit can be 0, 1, 2,... 9. Use this information and answer questions 3a and 3b
Question 3b: For your own safety, the bank representative says that the following choices are prohibited: 1) all four digits identical, 2) sequence of consecutive ascending or descending digits, such as 0123 or 3210, 3) any sequence starting with 19 or 20 (birth years). How many sequences are NOT permitted?
Answer #1

3b) not permitted sequence =N(all identical)+N(ascending )+N(descending)+N(starting with 19)+N(starting with 20)

=10+7+7+100+100 =224